# Grounded Reinforcement Learning for Visual Reasoning

### v1

@@ 1 Introduction
We hypothesize that models both “see better” and “think better” when their textual reasoning steps are explicitly grounded in specific image regions, promoting more targeted and systematic cross-referencing between textual and visual information during reasoning. This hypothesis is inspired by the fact that humans systematically shift their restricted gaze to selectively gather and integrate task-relevant information when reasoning about the world [75, 22, 74]. Grounding may serve a similar role in models, functioning as a spatial attention mechanism that enables accurate feature binding [61] and supports deictic reference [4] to simplify multi-step reasoning through localized perceptual anchoring. We move beyond prompt-based reasoning, proposing that learning to compose reasoning steps explicitly anchored in image coordinates induces structured region-level behaviors that support improved generalization in visual tasks.
@@ 3 Preliminaries
3.1 Do Current RL Recipes Amplify VLM Behaviors That Support Visual Performance?
@@ 4.3 Reinforcement Learning with Spatially-Grounded Reasoning Steps
Although $\pi_{\theta_{0}}$ imitates high-quality traces, it does not reason optimally for new queries. We therefore apply Group Relative Policy Optimization (GRPO; 50) to directly maximize task reward while preserving fluency and grounding (Figure 3, right). More details can be found in Appendix A5.1.
Reward Design. We wish to incentivize reasoning traces that include explicit grounding coordinates, and reasoning traces that lead to the final answer. Our total reward is a weighted sum: $R(\mathcal{T})=\lambda_{\text{fmt}}r_{\text{fmt}}+\lambda_{\text{task}}r_{\text{task}},$ where $r_{\text{fmt}}$ encourages valid and interpretable output format, and $r_{\text{task}}$ captures task-specific correctness. Importantly, along with checking outputs are formatted with correct <think> and <answer> tags, we award $+1$ $r_{\text{fmt}}$ only if all coordinate references are valid. We provide more details on task reward in Appendix A5.2.
@@ 4.3.1 Multi-Turn Reinforcement Learning for Visual Feedback
Grounded reasoning chains generated by our MCTS procedure already push the model to look at many regions, but the visual encoder still processes the same globally resized image at every step. Fine-grained cues (small text, icons, object boundaries) are therefore blurred away, potentially limiting the benefit of additional reasoning if image region details cannot be percieved by the model. Inspired by how humans zoom in after selecting a candidate region, we let the model request a higher-resolution crop, $o_{t}$, after predicting a coordinate. This “interactive microscope” supplies fresh evidence at a detail level impossible to encode in the initial global view.
@@ 5.1 RQ1: How much does grounded reasoning help on visual reasoning tasks?
1. Method Comparison: We compare ViGoRL to baselines utilizing the same training data and base model (Qwen-2.5-VL) as used in our method, but differing in their training recipe: 1) SFT-direct: Supervised fine-tuning on our trajectory dataset using final answers only. 2) Vanilla GRPO: GRPO applied to the base model with standard rewards for <think> and <answer> formatting, and answer correctness, similar to previous work [37, 38].
@@
Spatial Reasoning. We evaluate on SAT-2 [48] validation (4,000 questions across 5 categories), BLINK [11] (depth ordering, multi-view reasoning, spatial relationships), and ROBOSPATIAL-configuration and compatibility split (228 questions on real-world RGBD scenes) [53]. These benchmarks test generalization to novel environments, objects, and language configurations. Accuracy is measured via multiple-choice answer matching.
GUI Understanding. For grounding evaluation, we use ScreenSpot v2 [70, 6] (single-step localization), and ScreenSpot Pro [31] (high-resolution professional environments). To test small element grounding in low resolution, we additionally evaluate on ScreenSpot Pro with images downsampled to a lower resolution of 1920x1920, which we call ScreenSpot-Pro-LR (results in Appendix Table A6). Performance is measured by checking if predicted coordinates fall within target element bounding boxes. We also test on VisualWebArena [29] live web evaluation, using the visual-only configuration where models receive set-of-marks annotated webpage screenshots without HTML or text inputs. Task success is determined automatically by checking for specific criteria in the agent trajectory (see Koh et al. [29] for details).
Visual Search. The visual search benchmark VBench tests fine-grained visual understanding using 191 high-resolution images from the SA-1B dataset (avg. 2246×1582). It includes 115 attribute recognition samples (e.g., color, material) and 76 spatial relationship samples, evaluating models’ ability to analyze detailed object properties and relative positions in complex scenes.
@@ 5.1.2 Results
Qwen2.5-VL-3B 68.4% (±2.6) 9.3% (±1.4) 4.2% (±1.3)
+ SFT direct 80.6% (±2.2) 25.0% (±2.1) 4.5% (±1.4)
+ Vanilla GRPO 84.4% (±2.0) 29.0% (±2.2) 4.8% (±1.4)
ViGoRL-3b (Ours) 86.5% (±1.9) 31.1% (±2.3) 6.4% (±1.5)
ViGoRL-3b (Multi-t) 86.1% (±1.9) 32.3% (±2.4) -
Qwen2.5-VL-7B 73.6% (±2.4) 14.8% (±1.8) 5.5% (±1.5)
ViGoRL-7b (Ours) 91.0% (±1.6) 33.1% (±2.3) 11.2% (±2.1)
@@
Model SAT-2 Val BLINK RoboSpatial
@@
GPT-4o 60.0% 76.2%
GPT-4 Turbo 54.6%
Claude 3 Opus 44.1%
Gemini Pro 1.0 45.2%
@@
LLaVA v1.6 34B 46.8%
instructBLIP 13B 42.2%
Molmo - 67.1%
@@
Qwen2.5VL-3B 46.1% (±1.5) 44.4% (±2.3) 54.4% (±6.5)
+ SFT direct 58.3% (±1.5) 46.4% (±2.3) 62.3% (±6.3)
+ Vanilla GRPO 50.0% (±1.6) 46.5% (±2.3) 69.7% (±6.0)
ViGoRL-3b (Ours) 62.9% (±1.5) 48.5% (±2.3) 67.1% (±6.1)
Qwen2.5-VL-7B 52.6% (±1.6) 52.9% (±2.3) 46.7% (±9.5)
ViGoRL-7b (Ours) 67.5% (±1.5) 54.1% (±2.3) 76.4% (±7.5)
@@
Qwen2.5-7B-VL 78.0 (±5.9)
Multi-t ViGoRL-7B (Ours) 86.4 (±5.0)
Qwen2.5-3B-VL 74.2 (±6.2)
Multi-t ViGoRL-3B (Ours) 81.2 (±5.6)

### v2

@@ 1 Introduction
We hypothesize that models both “see better” and “think better” when their textual reasoning steps are explicitly grounded in specific image regions, promoting more targeted and systematic cross-referencing between textual and visual information during reasoning. This hypothesis is inspired by the fact that humans systematically shift their restricted gaze to selectively gather and integrate task-relevant information when reasoning about the world [77, 24, 76]. Grounding may serve a similar role in models, functioning as a spatial attention mechanism that enables accurate feature binding [63, 5, 6] and supports deictic reference [4] to simplify multi-step reasoning through localized perceptual anchoring. We move beyond prompt-based reasoning, proposing that learning to compose reasoning steps explicitly anchored in image coordinates induces structured region-level behaviors that support improved generalization in visual tasks.
@@ 3 Preliminaries
3.1 Do Current RL Recipes Amplify VLM Behaviors That Support Visual Reasoning?
@@ 4.3 Reinforcement Learning with Spatially-Grounded Reasoning Steps
Although $\pi_{\theta_{0}}$ imitates high‑quality traces, it does not reason optimally for new queries. We therefore apply Group Relative Policy Optimization (GRPO; 52) to directly maximize task reward while preserving fluency and grounding (Figure 3, right). More details can be found in Appendix A5.1.
Reward Design. We wish to incentivize reasoning traces that include explicit grounding coordinates, and reasoning traces that lead to the final answer. Our total reward is a weighted sum: $R(\mathcal{T})=\lambda_{\text{fmt}}r_{\text{fmt}}+\lambda_{\text{task}}r_{\text{task}},$ where $r_{\text{fmt}}$ encourages valid and interpretable output format, and $r_{\text{task}}$ captures task-specific correctness. Importantly, along with checking outputs are formatted with correct <think> and <answer> tags, we award +1+1 $r_{\text{fmt}}$ only if all coordinate references are valid. We provide more details on task reward in Appendix A5.2.
@@ 4.3.1 Multi-Turn Reinforcement Learning for Visual Feedback
Grounded reasoning chains generated by our MCTS procedure already push the model to look at many regions, but the visual encoder still processes the same globally resized image at every step. Fine-grained cues (small text, icons, object boundaries) are therefore blurred away, potentially limiting the benefit of additional reasoning if image region details cannot be perceived by the model. Inspired by how humans zoom in after selecting a candidate region, we let the model request a higher-resolution crop, $o_{t}$, after predicting a coordinate. This “interactive microscope” supplies fresh evidence at a detail level impossible to encode in the initial global view.
@@ 5.1 RQ1: How much does grounded reasoning help on visual reasoning tasks?
1. Method Comparison: We compare ViGoRL to baselines utilizing the same training data and base model (Qwen-2.5-VL) as used in our method, but differing in their training recipe: 1) SFT-direct: Supervised fine-tuning on our trajectory dataset using final answers only. 2) Vanilla GRPO: GRPO applied to the base model with standard rewards for <think> and <answer> formatting, and answer correctness, similar to previous work [39, 40]. Our model and all baselines decode with a temperature of 0.5 during evaluation.
@@
Spatial Reasoning. We evaluate on SAT-2 [50] validation (4,000 questions across 5 categories), BLINK [13] (depth ordering, multi-view reasoning, spatial relationships), and ROBOSPATIAL-configuration and compatibility split (228 questions on real-world RGBD scenes) [55]. These benchmarks test generalization to novel environments, objects, and language configurations. Accuracy is measured via multiple-choice answer matching. We use a max side length of 1260 pixels while keeping aspect ratio for evaluation.
GUI Understanding. For grounding evaluation, we use ScreenSpot v2 [72, 8] (single-step localization), and ScreenSpot Pro [33] (high-resolution professional environments). To test small element grounding in low resolution, we additionally evaluate on ScreenSpot Pro with images downsampled to a lower resolution of 1920x1920, which we call ScreenSpot-Pro-LR (results in Appendix Table A6). Performance is measured by checking if predicted coordinates fall within target element bounding boxes. We also test on VisualWebArena [31] live web evaluation, using the visual-only configuration where models receive set-of-marks annotated webpage screenshots without HTML or text inputs. Task success is determined automatically by checking for specific criteria in the agent trajectory (see Koh et al. [31] for details). We use a resize image to have maximum pixel size of 3600x3600.
Visual Search. The visual search benchmark VBench tests fine-grained visual understanding using 191 high-resolution images from the SA-1B dataset (avg. 2246×1582). It includes 115 attribute recognition samples (e.g., color, material) and 76 spatial relationship samples, evaluating models’ ability to analyze detailed object properties and relative positions in complex scenes. We use a resize image to have maximum pixel size of 3600x3600.
@@ 5.1.2 Results
Qwen2.5-VL-3B 68.4 (±2.6) 23.9 4.2 (±1.3)
+ SFT direct 80.6 (±2.2) 25.0 (±2.1) 4.5 (±1.4)
+ Vanilla GRPO 84.4 (±2.0) 29.0 (±2.2) 4.8 (±1.4)
ViGoRL-3b (Ours) 86.5 (±1.9) 31.1 (±2.3) 6.4 (±1.5)
Multi-t ViGoRL-3b (Ours) 86.1 (±1.9) 32.3 (±2.4) -
Qwen2.5-VL-7B 73.6 (±2.4) 29.0 5.5 (±1.5)
ViGoRL-7b (Ours) 91.0 (±1.6) 33.1 (±2.3) 11.2 (±2.1)
@@
Model SAT‑2 Val BLINK RoboSpatial
@@
GPT-4o 60.0 76.2
GPT-4 Turbo 54.6
Claude 3 Opus 44.1
Gemini Pro 1.0 45.2
@@
LLaVA v1.6 34B 46.8
instructBLIP 13B 42.2
Molmo - 67.1
@@
Qwen2.5VL-3B 46.1 (±1.5) 44.4 (±2.3) 54.4 (±6.5)
+ SFT direct 58.3 (±1.5) 46.4 (±2.3) 62.3 (±6.3)
+ Vanilla GRPO 50.0 (±1.6) 46.5 (±2.3) 69.7 (±6.0)
ViGoRL-3b (Ours) 62.9 (±1.5) 48.5 (±2.3) 67.1 (±6.1)
Qwen2.5-VL-7B 52.6 (±1.6) 52.9 (±2.3) 46.7 (±9.5)
ViGoRL-7b (Ours) 67.5 (±1.5) 54.1 (±2.3) 76.4 (±7.5)
@@
Qwen2.5-7B-VL 78.0 (±3.0)
Multi-t ViGoRL-7B 86.4 (±2.7)
Qwen2.5-3B-VL 74.2 (±3.1)
ViGoRL-3B (Ours) 79.1 (±3.0)
Multi-t ViGoRL-3B 81.2 (±2.8)
w/o diversity reward 78.0 (±3.0)
w/ bounding box outputs 81.2 (±2.8)
@@ 5.2 RQ2: Ablation Studies
6. Pointing is sufficient for visual search. As shown in Table 3, we observe no difference in V*Bench accuracy when switching to training ViGoRL to output variable size bounding box crops as opposed to the fixed-size point cropping used in ViGoRL. Fixed crops usually provide enough context and resolution for most cases, and ViGoRL can handle larger regions through multiple overlapping crops that together capture necessary information. Adaptive cropping could be more efficient but is limited by imprecise bounding boxes and inconsistent aspect ratios.